OpenStudy (anonymous):

How do I solve f(x)=x^2+1

OpenStudy (jdoe0001):

set f(x) to 0, then "solve" or get the x-intercepts, for "x" \(\bf f(x)=x^2+1\implies 0=x^2+1\)

OpenStudy (jdoe0001):

bear in mind that'd end up with a complex or "imaginary" result